OpenStudy (anonymous):

The train station clock runs too fast and gains 4 minutes every 3 days. How many minutes and seconds will it have gained at the end of 7 days?

OpenStudy (anonymous):

i did 4/3(7)= 28/3 or 9.33 so it would be 9 minutes and how many seconds? how do i figure out the seconds?

OpenStudy (turingtest):

first of all don't round it just yet... secondly, there are 60 seconds in every minute, so multiplying the number of minutes by the number of seconds in a minute gives the number of seconds.

OpenStudy (anonymous):

\[9\tfrac{1}{3}\]

OpenStudy (anonymous):

1/3 of a minute is 20 seconds if i am not mistaken

OpenStudy (anonymous):

how did you get 9 1/3?

OpenStudy (anonymous):

you said it \[\frac{28}{3}=9\tfrac{1}{3}\]